# LLM-FE: Automated Feature Engineering for Tabular Data with LLMs as Evolutionary Optimizers

### v1

@@ Abstract
Automated feature engineering plays a critical role in improving predictive model performance for tabular learning tasks. Traditional automated feature engineering methods are limited by their reliance on pre-defined transformations within fixed, manually designed search spaces, often neglecting domain knowledge. Recent advances using Large Language Models (LLMs) have enabled the integration of domain knowledge into the feature engineering process. However, existing LLM-based approaches use direct prompting or rely solely on validation scores for feature selection, failing to leverage insights from prior feature discovery experiments or establish meaningful reasoning between feature generation and data-driven performance. To address these challenges, we propose LLM-FE, a novel framework that combines evolutionary search with the domain knowledge and reasoning capabilities of LLMs to automatically discover effective features for tabular learning tasks. LLM-FE formulates feature engineering as a program search problem, where LLMs propose new feature transformation programs iteratively, and data-driven feedback guides the search process. Our results demonstrate that LLM-FE consistently outperforms state-of-the-art baselines, significantly enhancing the performance of tabular prediction models across diverse classification and regression benchmarks. [^1] The code and data are available at: https://github.com/nikhilsab/LLMFE
Machine Learning, ICML
@@ 1 Introduction
Feature engineering, the process of transforming raw data into meaningful features for machine learning models, is crucial for improving predictive performance, particularly when working with tabular data (Domingos, 2012). In many tabular prediction tasks, well-designed features have been shown to significantly enhance the performance of tree-based models, often outperforming deep learning models that rely on learned representations (Grinsztajn et al., 2022). However, data-centric tasks such as feature engineering are one of the most challenging processes in the tabular learning workflow (Anaconda, 2020; Hollmann et al., 2024), as they require experts and data scientists to explore many possible combinations in the vast combinatorial space of feature transformations. Classical feature engineering methods (Kanter & Veeramachaneni, 2015; Khurana et al., 2016, 2018; Horn et al., 2020; Zhang et al., 2023) construct extensive search spaces of feature processing operations, relying on various search and optimization techniques to identify the most effective features. However, these search spaces are mostly constrained by predefined, manually designed transformations and often fail to incorporate domain knowledge (Zhang et al., 2023). Domain knowledge can serve as an invaluable prior for identifying these transformations, leading to reduced complexity and more interpretable and effective features (Hollmann et al., 2024).
@@
Recently, Large Language Models (LLMs) have emerged as a powerful solution to this challenge, offering access to extensive embedded domain knowledge that can be leveraged for feature engineering. While recent approaches have demonstrated promising results in incorporating this knowledge into automated feature discovery, current LLM-based methods (Hollmann et al., 2024; Han et al., 2024) rely predominantly on direct prompting mechanisms or validation scores to guide the feature generation process. These approaches do not leverage insights from prior feature discovery experiments, thereby falling short of establishing meaningful reasoning between feature generation and data-driven performance.
@@
Figure 1: Overview of the LLM-FE Framework. For a given dataset, LLM-FE follows these steps: (a) New Hypothesis Generation, where an LLM generates feature transformation hypotheses as programs for a given tabular dataset; (b) Feature Engineering, where the feature transformation program is applied to the underlying dataset, resulting in a modified dataset; (c) Model Fitting, where a prediction model is fitted on the modified training set and evaluated on the corresponding validation set; (d) Multi-Population Memory, which maintains a buffer of high-scoring programs that act as in-context samples for LLM’s iterative refinement prompt.
@@
To address these limitations, we propose LLM-FE, a novel framework integrating the capabilities of LLMs with tabular prediction models and evolutionary search to facilitate effective feature optimization. As shown in Figure 1, LLM-FE follows an iterative process to generate and evaluate the hypothesis of the feature transformation, using the performance of the tabular prediction model as a reward to enhance the generation of effective features. Starting from an initial feature transformation program, LLM-FE leverages the LLMs’ embedded domain knowledge by incorporating task-specific details, feature descriptions, and a subset of data samples to generate new feature discovery hypotheses (Figure 1(a)). At each iteration, LLM acts as a knowledge-guided evolutionary optimizer, which mutates examples of previously successful feature transformation programs to generate new effective features (Meyerson et al., 2024). The refined hypotheses are then evaluated by augmenting the original dataset with the proposed feature hypotheses (Figure 1(b)) and training a tabular prediction model on the augmented data (Figure 1(c)). The model’s performance, measured on a held-out validation set, provides data-driven feedback that, combined with a dynamic memory of previously explored feature transformation programs (Figure 1(d)), guides the LLM to refine its hypothesis generation iteratively.
Table 1 summarizes a list of distinguishing features of LLM-FE compared to the state-of-the-art classical and LLM-based feature engineering methods. Unlike prior methods that rely on either fixed rules or pure LLM generation, LLM-FE leverages LLMs’ domain knowledge to seed a flexible hypothesis space while employing evolutionary optimization to iteratively refine features. This facilitates open-ended feature discovery, generalizing effectively across both classification and regression tasks in tabular data prediction.
@@
We evaluate LLM-FE with GPT-3.5-Turbo (OpenAI, 2023) and Llama-3.1-8B-Instruct (Dubey et al., 2024) backbones on classification and regression tasks across diverse tabular datasets. LLM-FE consistently outperforms the state-of-the-art feature engineering methods, identifying contextually relevant features that improve downstream performance. In particular, we observe improvements with tabular models like XGBoost (Chen & Guestrin, 2016), TabPFN (Hollmann et al., 2022), and MLP (Gorishniy et al., 2021). Our analysis also highlights the importance of evolutionary search in achieving effective results. The major contributions of this work can be summarized as follows.
@@
Table 1: Comparison of the proposed LLM-FE with existing feature engineering methods.
Method Classification Regression Domain Evolutionary
Knowledge Refinement
AutoFeat
OpenFE
FeatLLM
CAAFE
LLM-FE
@@ 2 Related Works
Feature engineering involves creating meaningful features from raw data to improve predictive performance (Hollmann et al., 2024). The growing complexity of datasets has driven the automation of feature engineering to reduce manual effort and optimize feature discovery. Traditional automated feature engineering methods include tree-based exploration (Khurana et al., 2016), iterative subsampling (Horn et al., 2020), and transformation enumeration (Kanter & Veeramachaneni, 2015). Learning-based methods leverage machine learning and reinforcement learning for feature transformation (Nargesian et al., 2017; Khurana et al., 2018; Zhang et al., 2019). OpenFE (Zhang et al., 2023) integrates a feature-boosting algorithm with a two-stage pruning strategy. These traditional approaches often fail to leverage domain knowledge for feature discovery, making LLMs well-suited for such tabular prediction tasks due to their prior contextual domain understanding.
@@
Recent advances in LLMs enable them to leverage their pre-trained knowledge to handle novel tasks through techniques such as prompt engineering and in-context learning without requiring additional training (Brown et al., 2020; Wei et al., 2022). Despite the progress made with LLMs, they often struggle with factually incorrect or inconsistent outputs (Madaan et al., 2024; Zhu et al., 2023). Researchers have thus explored methods that use feedback or refinement mechanisms to improve LLM responses and leverage them within complex optimization tasks (Madaan et al., 2024; Haluptzok et al., 2022). More recent approaches involving evolutionary optimization frameworks couple LLMs with evaluators (Lehman et al., 2023; Liu et al., 2024; Wu et al., 2024; Lange et al., 2024), using LLMs to perform adaptive mutation and crossover operations (Meyerson et al., 2024). This approach has shown success in areas such as prompt optimization (Yang et al., 2024; Guo et al., 2023), neural architecture search (Zheng et al., 2023; Chen et al., 2024), discovery of mathematical heuristics (Romera-Paredes et al., 2024), and symbolic regression (Shojaee et al., 2024). Building on these concepts, our LLM-FE framework utilizes an LLM as an evolutionary optimizer by coupling its prior knowledge with data-driven refinement to discover optimal features for the underlying tabular learning task.
@@
The rich prior knowledge encapsulated in LLMs has been harnessed to analyze structured data by serializing tabular data into natural language formats (Dinh et al., 2022; Hegselmann et al., 2023; Wang et al., 2023). Existing approaches include table-specific tokenization methods to pre-train models for consistent performance across diverse datasets (Yan et al., 2024) or employ fine-tuning/few-shot examples via in-context learning (Dinh et al., 2022; Hegselmann et al., 2023; Nam et al., 2023) to adapt LLMs to tabular prediction tasks. Recent research has explored the potential of LLMs for feature engineering. FeatLLM (Han et al., 2024) improves tabular predictions by generating and parsing rules to engineer binary features. CAAFE (Hollmann et al., 2024) introduces a context-aware approach where LLMs generate features directly from task descriptions, while OCTree (Nam et al., 2024) incorporates an additional decision tree reasoning feedback to enhance feature engineering. We advance this line of research by leveraging LLMs to efficiently navigate the optimization space of feature discovery, generating meaningful features that are informed by prior knowledge and enriched with data-driven insights and evolutionary exploration.
3 LLM-FE
@@ 3.1 Problem Formulation
A tabular dataset $\mathcal{D}$ comprises $N$ rows (or instances), each characterized by $d$ columns (or features). Each data instance $x_{i}$ is a $d$-dimensional feature vector with feature names denoted by $C=\{c_{j}\}_{j=1}^{d}$. The dataset is accompanied by metadata $\mathcal{M}$, which contains feature descriptions and task-specific information. For supervised learning tasks, each instance $x_{i}$ is associated with a corresponding label $y_{i}$, where $y_{i}\in\{0,1,...,K\}$ for classification tasks with $K$ classes, and $y_{i}\in\mathbb{R}$ for regression tasks. Given a labeled tabular dataset $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$, the primary objective is to derive a prediction model $f$ through the empirical risk minimization process. This model $f$ is designed to establish a mapping from the input feature space $\mathcal{X}$ to its corresponding label space $\mathcal{Y}$:
$\min_{f}\mathcal{L}_{f}(f(\mathcal{X}),\mathcal{Y})$ (1)
where $\mathcal{L}_{f}$ is the loss function. Our objective in feature engineering is to determine an optimal feature transformation $\mathcal{T}$, which enhances the performance of a predictive model when trained on the transformed input space. Formally, the feature engineering task can be defined as:
@@
subject to
$f^{*}=\operatorname*{arg\,min}_{f}\mathcal{L}_{f}(f(\mathcal{T}(\mathcal{X}_{\text{train}})),\mathcal{Y}_{\text{train}})$ (3)
where $\mathcal{T}$ is the feature transformation generated by the LLM $\pi_{\theta}$ and defined as $\mathcal{T}=\pi_{\theta}(\mathcal{X}_{train})$ meaning the transformation is learned from the training data by the LLM. The predictive model $f^{*}$ is then trained on the transformed training data $\mathcal{T}(\mathcal{X}_{train})$ to minimize loss. Consequently, the bilevel optimization problem seeks to identify the feature transformations $\mathcal{T}$ that maximize the performance $\mathcal{E}$ on the validation set $\mathcal{T}(\mathcal{X}_{\text{val}})$ while minimizing the loss function on the transformed training data, thereby efficiently exploring the potential feature space.
3.2 Hypothesis Generation
Figure 1(a) illustrates the hypothesis generation step that uses an LLM to create multiple new feature transformation programs, leveraging the model’s prior knowledge, reasoning, and in-context learning abilities to effectively explore the feature space.
@@ 3.2.1 Input Prompt
The LLM is assigned the task of finding the most relevant features to help solve the given regression/ classification problem. The task emphasizes using the LLM’s prior knowledge of the dataset’s domain to generate features. The LLM is explicitly instructed to generate novel features and provide clear step-by-step reasoning for their relevance to the prediction task.
@@
After providing the instructions, we provide LLM with the dataset-specific information from the metadata $\mathcal{M}$. This information encompasses a detailed description of the intended downstream task, along with the feature names $C$ and their corresponding descriptions. In addition, we provide a limited number of representative samples from the tabular dataset. To improve the effective interpretation of the data, we adopt the serialization approach used in previous works (Dinh et al., 2022; Hegselmann et al., 2023; Han et al., 2024). We serialized the data samples as follows:
$\small\text{Serialize}(x_{i},y_{i},C)=$
`$\text{If }c_{1}\text{ is }x_{i}^{1},...,c_{d}\text{ is }x_{i}^{d}.~{}~{}\text{Then Result is }y_{i}\text{'}$ ’ (4)
@@
3.2.2 Hypothesis Sampling
At each iteration $t$, we construct the prompt $p_{t}$ by sampling the previous iteration as input to the LLM $\pi_{\theta}$, resulting in the output $\mathcal{T}_{1},\dots,\mathcal{T}_{b}=\pi_{\theta}(p_{t})$ representing a set of $b$ sampled programs. To promote diversity and maintain a balance between exploration (creativity) and exploitation (prior knowledge), we employ stochastic temperature-based sampling. Each of the sampled feature hypotheses ($\mathcal{T}_{i}$) is executed before evaluation to discard error-prone programs. This ensures that only valid and executable feature transformation programs are considered further in the optimization pipeline. In addition, to ensure computational efficiency, a maximum execution time threshold is enforced, discarding any programs that exceed it.

### v2

@@ Abstract
Automated feature engineering plays a critical role in improving predictive model performance for tabular learning tasks. Traditional automated feature engineering methods are limited by their reliance on pre-defined transformations within fixed, manually designed search spaces, often neglecting domain knowledge. Recent advances using Large Language Models (LLMs) have enabled the integration of domain knowledge into the feature engineering process. However, existing LLM-based approaches use direct prompting or rely solely on validation scores for feature selection, failing to leverage insights from prior feature discovery experiments or establish meaningful reasoning between feature generation and data-driven performance. To address these challenges, we propose LLM-FE, a novel framework that combines evolutionary search with the domain knowledge and reasoning capabilities of LLMs to automatically discover effective features for tabular learning tasks. LLM-FE formulates feature engineering as a program search problem, where LLMs propose new feature transformation programs iteratively, and data-driven feedback guides the search process. Our results demonstrate that LLM-FE consistently outperforms state-of-the-art baselines, significantly enhancing the performance of tabular prediction models across diverse classification and regression benchmarks.
The code is available at: [Uncaptioned image]:https://github.com/nikhilsab/LLMFE
@@ 1 Introduction
Feature engineering, the process of transforming raw data into meaningful features for machine learning models, is crucial for improving predictive performance, particularly when working with tabular data domingos2012few . In many tabular prediction tasks, well-designed features have been shown to significantly enhance the performance of tree-based models, often outperforming deep learning models that rely on learned representations grinsztajn2022tree . However, data-centric tasks such as feature engineering are one of the most challenging processes in the tabular learning workflow anaconda2020state ; hollmann2024large , as they require experts and data scientists to explore many possible combinations in the vast combinatorial space of feature transformations. Classical feature engineering methods kanter2015deep ; khurana2016cognito ; khurana2018feature ; horn2020autofeat ; zhang2023openfe construct extensive search spaces of feature processing operations, relying on various search and optimization techniques to identify the most effective features. However, these search spaces are mostly constrained by predefined, manually designed transformations and often fail to incorporate domain knowledge zhang2023openfe . Domain knowledge can serve as an invaluable prior for identifying these transformations, leading to reduced complexity and more interpretable and effective features hollmann2024large .
@@
Recently, Large Language Models (LLMs) have emerged as a powerful solution to this challenge, offering access to extensive embedded domain knowledge that can be leveraged for feature engineering. While recent approaches have demonstrated promising results in incorporating this knowledge into automated feature discovery, current LLM-based methods hollmann2024large ; han2024large rely predominantly on direct prompting mechanisms or validation scores to guide the feature generation process. These approaches do not leverage insights from prior feature discovery experiments, thereby falling short of establishing meaningful reasoning between feature generation and data-driven performance.
@@
Figure 1: Overview of the LLM-FE Framework. For a given dataset, LLM-FE follows these steps: (a) New Feature Generation, where an LLM generates feature transformation hypotheses as programs for a given tabular dataset; (b) Feature Engineering, where the feature transformation program is applied to the underlying dataset, resulting in a modified dataset; (c) Feature Evaluation, where the modified dataset with the new features is evaluated using a prediction model; (d) Experience Management, which maintains a buffer of high-scoring programs that act as in-context samples for LLM’s iterative refinement prompt. The features generated by LLM-FE are interpretable, using LLM’s knowledge of the problem’s domain.
@@
To address these limitations, we propose LLM-FE, a novel framework integrating the capabilities of LLMs with tabular prediction models and evolutionary search to facilitate effective feature optimization. As shown in Figure 1, LLM-FE follows an iterative process to generate and evaluate the hypothesis of the feature transformation, using the performance of the tabular prediction model as a reward to enhance the generation of effective features. Starting from an initial feature transformation program, LLM-FE leverages the LLMs’ embedded domain knowledge by incorporating task-specific details, feature descriptions, and a subset of data samples to generate new feature discovery programs (Figure 1(a)). At each iteration, LLM acts as a knowledge-guided evolutionary optimizer, which mutates examples of previously successful feature transformation programs to generate new effective features meyerson2024language . The newly proposed features are then integrated with the original dataset to yield an augmented dataset (Figure 1(b)). The prediction model’s performance is evaluated on a held-out validation set derived from the augmented dataset (Figure 1(c)), provides data-driven feedback that, combined with a dynamic memory of previously explored feature transformation programs (Figure 1(d)), guides the LLM to refine its feature generation iteratively.
Table 1: Comparison of existing feature engineering methods.
Method Domain Feedback Complex Multi-Feature
Knowledge Driven Features Refinement
AutoFeat horn2020autofeat
OpenFE zhang2023openfe
FeatLLM han2024large
CAAFE hollmann2024large
OCTree nam2024optimized
LLM-FE
@@
Table 1 compares LLM-FE to several state-of-the-art classical and LLM-based feature engineering methods. Traditional methods lack adaptability and deeper contextual understanding, while LLM-based methods generate simple features kuken2024large or use feedback to iteratively refine only a single rule. In contrast, LLM-FE supports all four aspects by leveraging LLM-based domain knowledge and feedback-driven optimization to generalize well across table prediction tasks. We evaluate LLM-FE with GPT-3.5-Turbo openai2023gpt and Llama-3.1-8B-Instruct dubey2024llama backbones on classification and regression tasks across diverse tabular datasets. LLM-FE consistently outperforms the state-of-the-art feature engineering methods, identifying contextually relevant features that improve downstream performance. In particular, we observe improvements with tabular models like XGBoost chen2016xgboost , TabPFN hollmann2022tabpfn , and MLP gorishniy2021revisiting . Our analysis also highlights the importance of evolutionary search in achieving effective results. The major contributions of this work can be summarized as.
@@ 2 Related Works
Feature engineering involves creating meaningful features from raw data to improve predictive performance hollmann2024large . The growing complexity of datasets has driven the automation of feature engineering to reduce manual effort and optimize feature discovery. Traditional automated feature engineering methods include tree-based exploration khurana2016cognito , iterative subsampling horn2020autofeat , and transformation enumeration kanter2015deep . Learning-based methods leverage machine learning and reinforcement learning for feature transformation nargesian2017learning ; khurana2018feature ; zhang2019automatic . OpenFE zhang2023openfe integrates a feature-boosting algorithm with a two-stage pruning strategy. These traditional approaches often fail to leverage domain knowledge for feature discovery, making LLMs well-suited for such tabular prediction tasks due to their prior contextual domain understanding.
@@
Recent advances in LLMs enable them to leverage their pre-trained knowledge to handle novel tasks through techniques such as prompt engineering and in-context learning without requiring additional training brown2020language ; wei2022chain . Despite the progress made with LLMs, they often struggle with factually incorrect or inconsistent outputs madaan2024self ; zhu2023large . Researchers have thus explored methods that use feedback or refinement mechanisms to improve LLM responses and leverage them within complex optimization tasks madaan2024self ; haluptzok2022language . More recent approaches involving evolutionary optimization frameworks couple LLMs with evaluators lehman2023evolution ; liu2024large ; wu2024evolutionary ; lange2024large , using LLMs to perform adaptive mutation and crossover operations meyerson2024language . This approach has shown success in areas such as prompt optimization yang2024largelanguagemodelsoptimizers ; guo2023connecting , neural architecture search zheng2023can ; chen2024evoprompting , discovery of mathematical heuristics romera2024mathematical , and symbolic regression shojaee2024llm . Building on these concepts, our LLM-FE framework utilizes an LLM as an evolutionary optimizer by coupling its prior knowledge with data-driven refinement to discover optimal features for the underlying tabular learning task.
@@
LLMs have been applied to structured data by converting tables into natural language dinh2022lift ; hegselmann2023tabllm ; wang2023anypredict . Approaches include table-specific tokenization for robust pre-training yan2024making and using fine-tuning or few-shot in-context learning for tabular prediction tasks dinh2022lift ; hegselmann2023tabllm ; nam2023stunt . Recently, LLMs have been used for feature engineering with FeatLLM han2024large , improving tabular predictions by generating and parsing rules to engineer binary features. CAAFE hollmann2024large introduces a context-aware approach where LLMs generate features directly from task descriptions, while OCTree nam2024optimized incorporates an additional decision tree reasoning step to iteratively improve a single rule at a time. In contrast, LLM-FE maintains multiple promising candidates and uses evolutionary search to efficiently navigate the feature space via mutation and crossover without using any additional steps. LLM-FE thus uses LLMs to efficiently navigate the optimization space of feature discovery, generating meaningful features that are informed by prior knowledge and enriched with data-driven insights and evolutionary exploration.
3 LLM-FE Approach
@@ 3.1 Problem Formulation
A tabular dataset $\mathcal{D}$ comprises $N$ rows (or instances), each characterized by $d$ columns (or features). Each data instance $x_{i}$ is a $d$-dimensional feature vector with feature names denoted by $C=\{c_{j}\}_{j=1}^{d}$. The dataset is accompanied by metadata $\mathcal{M}$, which contains feature descriptions and task-specific information. For supervised learning tasks, each instance $x_{i}$ is associated with a corresponding label $y_{i}$, where $y_{i}\in\{0,1,...,K\}$ for classification tasks with $K$ classes, and $y_{i}\in\mathbb{R}$ for regression tasks. Given a labeled tabular dataset $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$ and prediction model $f$ to map from the input feature space $\mathcal{X}$ to its corresponding label space $\mathcal{Y}$, the feature engineering objective is to determine an optimal feature transformation $\mathcal{T}$, which enhances the performance of a predictive model when trained on the transformed input space. Formally, the feature engineering task can be defined as:
@@
subject to:
$f^{*}=\operatorname*{arg\,min}_{f}\mathcal{L}_{f}(f(\mathcal{T}(\mathcal{X}_{\text{tr}})),\mathcal{Y}_{\text{tr}})$ (2)
where $(\mathcal{X}_{tr},\mathcal{Y}_{tr})$ and $(\mathcal{X}_{val},\mathcal{Y}_{val})$ are the sub-training set and validation set, respectively, that is derived from the training data $(\mathcal{X}_{train},\mathcal{Y}_{train})$. The feature transformation $\mathcal{T}$ generated by the LLM $\pi_{\theta}$ and defined as $\mathcal{T}=\pi_{\theta}(\mathcal{X}_{train})$, meaning the transformation is learned from the training data by the LLM. The predictive model $f^{*}$ is then trained on the transformed training data $\mathcal{T}(\mathcal{X}_{train})$ to minimize loss. Consequently, the bilevel optimization problem seeks to identify the feature transformations $\mathcal{T}$ that maximize the performance $\mathcal{E}$ on $\mathcal{T}(\mathcal{X}_{\text{val}})$ while minimizing the loss function on the transformed training data, thereby efficiently exploring the potential feature space.
3.2 Feature Generation
Figure 1(a) illustrates the feature generation step that uses an LLM to create multiple new feature transformation programs, leveraging the model’s prior knowledge, reasoning, and in-context learning abilities to effectively explore the feature space.
@@ 3.2.1 Input Prompt
The LLM is assigned the task of finding the most relevant features to help solve the given task. The task emphasizes using the LLM’s prior knowledge of the dataset’s domain to generate features. The LLM is explicitly instructed to generate novel features and provide clear step-by-step reasoning for their relevance to the prediction task. Moreover, since LLMs tend to generate simple features, we specifically instruct the LLM to generate complex features.
@@
After providing the instructions, we provide LLM with the dataset-specific information from the metadata $\mathcal{M}$. This information encompasses a detailed description of the intended downstream task, along with the feature names $C$ and their corresponding descriptions. In addition, we provide a limited number of representative samples from the tabular dataset. To improve the effective interpretation of the data, we adopt the serialization approach used in previous works dinh2022lift ; hegselmann2023tabllm ; han2024large . We serialized the data samples as follows:
$\text{Serialize}(x_{i},y_{i},C)=`\text{If }c_{1}\text{ is }x_{i}^{1},...,c_{d}\text{ is }x_{i}^{d}.\leavevmode\nobreak\ \leavevmode\nobreak\ \text{Then Result is }y_{i}\text{'}$ ’ (3)
@@
3.2.2 Feature Sampling
At each iteration $t$, we construct the prompt $p_{t}$ by sampling the previous iteration as input to the LLM $\pi_{\theta}$, resulting in the output $\mathcal{T}_{1},\dots,\mathcal{T}_{b}=\pi_{\theta}(p_{t})$ representing a set of $b$ sampled programs. To promote diversity and maintain a balance between exploration (creativity) and exploitation (prior knowledge), we employ stochastic temperature-based sampling. Each of the sampled feature transforms ($\mathcal{T}_{i}$) is executed before evaluation to discard error-prone programs. This ensures that only valid and executable feature transformation programs are considered further in the optimization pipeline. In addition, to ensure computational efficiency, a maximum execution time threshold is enforced, discarding any programs that exceed it.